OpenStudy (peterg2001):

please help will medal In order to receive credit, you must SHOW ALL YOUR WORK that is necessary to answer the question. Question 1: If f(x) = x2 - x + 2, then find the value for f(-3). Question 2: Evaluate 2xy - m when x = -4, y = 3, and m = 6 Question 3: Solve: -5 + x/3 = 20

OpenStudy (peterg2001):

@DanJS

OpenStudy (danjs):

hi

OpenStudy (danjs):

Which part do you not understand?

OpenStudy (danjs):

I'll get you started... Question 1) f(x) = x^2 - x + 2 f(-3)=? f(-3) = (-3)^2 - (-3) +2 f(-3) = 9 + 3 + 2 You just replace the variable with the numerical value, in each expression.

OpenStudy (crissy15):

question 2 is 2(-4)(3)-6 question 3 is i'm stumped how to get you started

OpenStudy (peterg2001):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

1. substitute -3 for x

OpenStudy (alexandervonhumboldt2):

9+3+2 for 1.

OpenStudy (alexandervonhumboldt2):

f(-x)=14

OpenStudy (peterg2001):

@abram134

OpenStudy (anonymous):

don't get it

OpenStudy (danjs):

Did you still need help here?

OpenStudy (danjs):

Evaluate 2xy - m when x = -4, y = 3, and m = 6 Replace each variable with the numbers 2xy - m = 2*(-4)*(3) - 6 = -30

OpenStudy (danjs):

Solve: -5 + x/3 = 20 Multiply through by 3 first to clear fractions -15 + x = 60 add 15 to both sides x = 75

OpenStudy (peterg2001):

@danJS thanks